OpenStudy (anonymous):

what is one of the roots of (z-9)^2=-4

OpenStudy (anonymous):

z=9+2i

OpenStudy (anonymous):

can you explain that to me please? :)

OpenStudy (anonymous):

z-9 =sqrt(-4) z= 9 +sqrt(-4) z= 9 +sqrt(4)*sqrt(-1) z= 9 +2*i

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

and as the roots appear as conjugate so the other root will be 9 - 2i